OpenStudy (anonymous):

how do you graph this? (x+2)^2+(y+1)^2=121 I simplified it to x^2+y^2=116 but I don't know if I was supposed to do that... HELP!

jimthompson5910 (jim_thompson5910):

(x+2)^2+(y+1)^2=121 (x-(-2))^2+(y-(-1))^2=11^2 so this is a circle with center: (-2, -1) radius: 11

OpenStudy (anonymous):

 So, I'm no artist, but does it look like this?